The following table shows the Myers-Briggs personality preferences for a random sample of 400 people in the listed professions.

Occupation Extroverted Introverted Row Total
Clergy (all denominations) 61 46 107
M.D. 68 90 158
Lawyer 53 82 135
Column Total 182 218 400

Use the chi-square test to determine if the listed occupations and personality preferences are independent at the 0.1 level of significance. What are the degrees of freedom?

Select one:

a. 5

b. 6

c. 2

d. 4

e. 3
Answer #1

Solution:

We know that the degrees of freedom of chi square test is given by ,

Df= (r-1)×(c-1)

Where , r = number of rows in a data and

c= number of columns in a data

Here , r= 3 and c = 2

Df = (3-1)×(2-1)

=2×1

= 2

Answer : c) 2